I'm trying to go through a systems of equations problem using elimination.
-2x+y=-2
2x+y=6

Answer by Drat1856(2)

(-2x+y=-2)
-(2x+y=6)
-2x-2x+y-y=-2-6
-4x=-8
-4x/-4=-8/-4
x=2
-2(2)+y=-2
-4+y=-2
-4+4+y=-2+4
y=2
2(2)+y=6
4+y=6
4-4+y=6-4
y=2
Therefore: x=2; y=2